OpenStudy (anonymous):

The dollar amount of United States exports to Russia increased from $1480 million to $2768 million between 1988 and 1999. What was the approximate percent increase? A. 83% B. 84% C. 85% D. 87%

OpenStudy (janu16):

$2,768 - $1,480 ___________ $1,288 So... ($1,288)/($1,480) = .8702702703 Meaning: the amount increased was .8702702703 (87%) of the original. so 87

OpenStudy (anonymous):

thankyou

OpenStudy (mahoganie.carson):

Welcome to Open Study $2768-1480=1288 1288/1480=87% @Swagmadness251

OpenStudy (anonymous):

thankyou